OpenStudy (anonymous):

Evaluate: 16 − 10 / 2 + 3 Evaluate: 16 − 10 / 2 + 3 @Mathematics

OpenStudy (anonymous):

8..

OpenStudy (anonymous):

14

OpenStudy (anonymous):

16-5+3=11+3=14

OpenStudy (anonymous):

i feel it is 29..........

OpenStudy (anonymous):

why again +3 there is one 3 in there

OpenStudy (anonymous):

no it sud be 45